Using the formula pH = -log[H+], show me how to find the hydrogen ion content of coffee when the pH level is 5.0
Write answer in scientific notation including units.

Answer by edjones(7311)

-log[H+]=5
10^(log(1/[H+]))=1/10^5
1/H+=1/10^5
.
Ed